One mole of copper(II) Nitrate, Cu(NO3)2 contains how many O atoms? (Avagadros number)
1
Expert's answer
2019-04-18T05:04:50-0400

Each molecule of copper (II) nitrate contains 6 atoms of Oxygen. One mole of any substance contains 6.02*1023 structural units. So, there will be 6 times more Oxygen atoms, than we have copper (II) nitrate molecules:


"N(O)=6*6.02*10^{23}=36.12*10^{23}=3.61210^{24} atoms"

where N - number of atoms.

Answer: one mole of copper (II) nitrate contains 3.612*1024 Oxygen atoms.